OpenStudy (anonymous):

When solving the system of equations y equals x squared, y equals 10 minus three x algebraically, which of the following quadratic equations do you need to solve during the process?

OpenStudy (anonymous):

x2 − 3x − 10 = 0 x2 + 3x − 10 = 0 x2 + 3x + 10 = 0 2y = x2 − 3x + 10

OpenStudy (anonymous):

Have any ideas?

OpenStudy (anonymous):

I think its the last one but not quite sure

OpenStudy (anonymous):

That would be incorrect because remember to solve a quadratic it's better if you have all terms on one side

OpenStudy (anonymous):

So what would this system be if you tried to solve it with substitution?

OpenStudy (anonymous):

maybe the first no clue honestly

OpenStudy (anonymous):

Ok so y=x^2 and y=10-3x So therefore since y=x^2 we can say \[x^2=10-3x\]

OpenStudy (anonymous):

Now try and get all the terms on 1 side